Activity

Optimistic(ta)

Most Likely(tm)

Pessimistic(tb)

Predecessor

A

4

5

6

             ___

B

5

6

8

             ___

C

3

4

7

             ___

D

4

5

6

             A

E

3

3

3

             A

F

3

4

8

             B

G

5

8

12

             C

H

1

1

1

             D

I

1

2

3

             E, F

J

2

3

6

             G

K

4

7

9

             I, J

L

2

5

7

             H, K

a. Find the following:

-Expected Time(te)

-Variance(σ2)

-ES, LS, EF, LF, and S.

-critical path.

b. Find the probability that the project will not be completed in the expected time (CT) + 5 days?
Answer #1

a)

Critical path is the longest path which is C-G-J-K-L having duration 27.5 days

Formula

b)

Expected completion time = duration of critical path = 27.5 days

z = (x-Expected completion time)/project standard deviation

= 5/sqrt(3.638888889)
= 2.62111217

probability of getting the project completed within expected time (CT) + 5 days = NORM.S.DIST(2.62111217,TRUE) = 0.995617829

So, probability that the project will not be completed in the expected time (CT) + 5 days = 1-0.995617829 = 0.004382171